TWO-COLOR, LASER-INDUCED GRATING SPECTROSCOPY: A MULTIDIMENSIONAL PROBE OF EXCITED STATE DYNAMICS

\underline{LEANNE M. MILLER}, THOMAS J. DOWD, and JAMES S. KELLER, Department of Chemistry and Biochemistry, University of Notre Dame, Notre Dame IN, 46556.

The population of vibrationally excited levels prior to photolysis alters the excited state dissociation dynamics through nuclear motion that is initiated on a ground electronic surface. Accordingly, we perform vibrationally mediated photodissocation to study novel excited state geometries and reaction pathways of small gas phase molecules. In a technique known as two-color laser-induced grating spectroscopy (TC-LIGS), a grating is formed by the interference of two degenerate “pump” beams tuned to an overtone of a light atom oscillator such as a O–H, N–H, or C–H bond. A third photolysis beam crosses the interference region and is diffracted from this grating, giving rise to a zero-background signal beam. Comparison of photoacoustic (one photon) and two-color grating (two photon) spectra, reveals the influence vibrational excitation has on the behavior of an electronically excited molecule.